# Low-Light Video Enhancement via Spatial-Temporal Consistent Illumination and Reflection Decomposition

### v1

@@ 4.5 User Study
To prove the effectiveness of our proposed framework in terms of human subjective evaluation, we conduct a large-scale user study with 100 participants with varying ages and education backgrounds, and balanced sex distribution. We aim to compare the perceptual quality of our method against SOTA baselines. To conduct the user study, we randomly collect 30 videos in the wild with the capture device of Sony A7R3 and iPhone13 (all models in the user study are trained with DID, since it is a large-scale dataset with multiple scenes and devices). Following most of the existing low-light enhancement works Xu et al. (2023b); Wang et al. (2023a), we employ the AB-test for the user study. For a video, the result produced by our method is called “Video A” whereas the result from another baseline is “Video B”. During the evaluation, each participant will compare the results of our method against other baselines in random order. For the comparison with one baseline, 5 videos will be randomly selected for visual perception. Each participant would simultaneously watch videos A and B. To avoid the bias caused by the position, we will randomize the left-right presentation order when showing videos A and B in each question. The participants can choose among three options: “Video A is better”, “Video B is better”, and “I think they are of the same quality”. They should make the decision according to the natural brightness, contrast, and color of each frame; the rich details, fewer artifacts, and the temporal consistency in the video. For each participant, the number of tasks is 10 methods $\times$ 5 videos $=50$, and it took around 50 minutes on average for each participant to complete the user study.
@@
Table 4: Quantitative comparison on 3D low-light datasets with normal-light pairs, proposed in Wang et al. (2023a).
Methods PairLLE SNR RetinexFormer SCI+NeRF Ours
PSNR 18.35 18.87 19.49 13.08 21.09
SSIM 0.604 0.650 0.702 0.505 0.763
Methods PairLLE+NeRF SNR+NeRF RetinexFormer+NeRF LLNeRF Ours+NeRF
PSNR 18.72 19.24 19.98 20.50 22.35
SSIM 0.681 0.693 0.736 0.758 0.782
4.6 Enhancement for NeRF
Recent works have introduced the implementation of 3D low-light enhancement Wang et al. (2023a), with a primary focus on disentangling view-independent intrinsic color and view-varying illuminations to achieve 3D-consistent enhancement. On the other hand, our framework’s enhanced results are particularly well-suited for NeRF reconstruction, as we have implemented a decomposition that maintains consistency across different spatial locations within a 3D scene.
To assess the impact of using our enhanced results for NeRF, we conducted experiments on a multi-view low-light dataset previously introduced in Wang et al. (2023a). We employed the network trained on DID, as it benefits from a large-scale training dataset, resulting in solid generalization capabilities. Furthermore, DID inherently contains multi-view images. Specifically, for each view, we selected its closest view as the reference view for our input data.
In Table 4, we present the PSNR and SSIM results on the testing set of Wang et al. (2023a), where ground truth data is available. It’s important to note that none of the networks were fine-tuned on the testing set. The superiority of our framework is evident, as our network can achieve 3D-consistent enhancement that aligns with the fundamental principles of NeRF.
@@ 5 Conclusion
We introduce a novel LLVE framework that utilizes spatial-temporal consistent illumination and reflection decomposition and dual networks for mutual feature propagation. It predicts the decomposition of normal-light outputs without external priors, thanks to cross-frame correspondences and practical illumination continuity. Extensive experiments on various datasets showcase the framework’s superiority over SOTA approaches and the impact of our designed components.

### v2

@@ 4.5 User Study
To prove the effectiveness of our proposed framework in terms of human subjective evaluation, we conduct a large-scale user study with 100 participants with varying ages and education backgrounds, and balanced sex distribution. Following most of the existing low-light enhancement works (Xu et al., 2023b; Wang et al., 2023a), we employ the AB-test for the user study. Participants should indicate their preference or select the same option. They should make the decision according to the natural brightness, contrast, and color of each frame; the rich details, fewer artifacts, and the temporal consistency in the video.
@@
4.6 The Effects of CFIM Structure
As previously discussed, the use of CFIM yields two key benefits: it enhances the consistency of the image enhancement process and strengthens the robustness of the backbone by providing the cross-attention mechanism with information from random neighboring time steps. In this section, we conduct experiments to show the effects of CFIM. The baseline consists of a network that processes a single frame, while the comparative setup integrates CFIM, where cross-attention is applied between randomly selected pairs of frames during training. The results of this comparison are presented in Table 5, which show that models incorporating the CFIM structure significantly outperform those without it, thereby demonstrating the effectiveness of the CFIM approach. Besides, the ablation results of the mentioned “w/o C.F.” (in Sec. 4.4) also support the effects of CFIM in the decomposition.
SDSD-indoor SDSD-outdoor SMID DID
Methods PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
SNR 27.30 0.84 23.23 0.82 28.49 0.81 24.85 0.90
SNR+CFIM 28.15 0.86 24.57 0.83 29.24 0.82 26.03 0.91
RetinexFormer 26.56 0.79 22.80 0.77 29.15 0.82 25.40 0.89
RetinexFormer+CFIM 27.42 0.80 23.75 0.79 29.87 0.83 26.38 0.90
Table 5: The experiments to show the effects of CFIM.
4.7 Robustness Towards Correspondences
In this section, we explore the robustness of our framework with respect to incorrect and insufficient correspondences. In the first experiment, we introduce random noise perturbations to the computed correspondences, modifying the location values of the correspondences (the perturbation range is sampled from $-20\sim 20$ pixels, which falls within the normal error range of current correspondence estimation methods (Edstedt et al., 2023)). As shown in Table 6, the results with perturbed correspondences (denoted as Ours with per.) show a performance drop compared to the unperturbed case. However, the performance remains superior to that of most SOTA baselines, indicating that our framework is robust to a certain of erroneous correspondences, which are likely to occur in areas with complex textures or other challenging scenarios. In the second experiment, we address the issue of insufficient correspondences, which can arise, for instance, in regions with occlusions. To simulate this, we randomly reduce the number of correspondences to 10% of the originally detected ones. The results (Ours with reduced) are presented in Table 6, where we observe that our method continues to outperform almost all baselines in Table 1, albeit with a smaller margin than the original setting. This suggests that even with fewer correspondences, the remaining ones still provide valuable information and they are in accord with the other constraints. This further demonstrates the robustness of our method.
SDSD-indoor SDSD-outdoor SMID DID
Methods PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
Ours with reduce 27.97 0.84 25.73 0.80 29.12 0.79 28.86 0.89
Ours with per. 28.10 0.86 26.04 0.79 29.37 0.81 29.08 0.91
Ours 28.93 0.88 26.32 0.82 29.60 0.82 30.10 0.93
Table 6: The experiments to show the robustness of our framework towards correspondences.
@@ 5 Conclusion
We introduce a novel LLVE framework that utilizes spatial-temporal consistent decomposition and dual networks for mutual feature propagation. It predicts the decomposition of normal-light outputs without external priors, thanks to cross-frame correspondences and practical constraint of the continuity. Extensive experiments on various datasets showcase the framework’s superiority over SOTA approaches and the impact of our designed components.
Ethical Statement
There are no ethical issues.
Acknowledgments
This work is supported by the Natural Science Foundation of Zhejiang Province, China, under No. LD24F020002.